The Emmett Bath Tub Company sells bathtub in West Virginia. The number sold for the past several months is given along with the forecasts using both 3-month moving average and exponential smoothing (α = .2). Determine the better forecast using the MSE.

Month   

Sales

3-month moving average

Error

Exponential   

Smoothing

Error   

Jan

13

Feb

18

March

15

Apr

12

May

14

June

15

July

10
Answer #1

Moving Average:

period 3 mont MA forecast error ABS(error) month Jan sales 13 18 Feb WN Mar Apr May Jun 15.33 15.00 13.67 13.67 -1.00 1.33 -3

Formula Ref:

period sales 3 mont MA forecast error ABS(error) month Jan Feb Mar NO Apr May =AVERAGE(C6:08) =AVERAGE(C7:09) =AVERAGE(C8:C10

Exponential:

Alpha = 0.2 period sales error ABS(error) month Jan Feb Mar Apr May Jun forecast 13.00 13.00 14.00 14.20 13.76 13.81 14.05 5.

Formula Ref:

Alpha = 0.2 period sales ABS(error) 4 5 1 62 month Jan Feb Mar 8 4 forecast error =C5 =$D$3*C5+(1-$D$3)*D5 =C6-D6 =$D$3*C6+(1